find the slope of th line parallel to the line 2x + 3y = -3
Answer by stanbon(75887)

find the slope of th line parallel to the line 2x + 3y = -3
------------
Put the equation in slope-intercept form:
3y = -2x -3
y = (-2/3)x + 1
=================
Slope = -2/3 ; intercept = 1
====================
Cheers,
Stan H.